find the roots of the following problems:
5squareroot of x,-x=6.


and x^2-8=17.

show all work

Answer by RAY100(1637)

5sqrtx,,,,,with -x=6 or x=-6
.
5 sqrt (-6)
.
5i(sqrt6),,,,,,with i=sqrt(-1)
.
.
or if problem is ,,,,,,5sqrtx -x =6
.
5sqrtx =x+6,,,,,,,,square both sides
.
25*x =(x+6)^2 = x^2 +12x +36
.
0= x^2 -13x +36,,,,,factor
.
(x-9)(x-4) =0
.
x=9,,,,4
.
.
2nd problem,,,,,x^2 -8 = 17
.
x^2 = 9
.
x= +/- 3
.
.